Reports 1-1 of 1 Clear search Modify search
MIF (ITF Control)
takafumi.ushiba - 22:51 Friday 28 October 2022 (22726) Print this report
PRMI lock trial Day2: PRMI can be locked with 3F signals

Continued work of klog22711

PRMI can be locked with 3F signals (REFL 51Q for MICH and REFL 135I for PRCL) today.
So, I designed the control filters and measure the OLTFs of MICH and PRCL.


Today, I discussed with Miyakawa-san on the initial alignment procedure for PRC and concluded that it is better to use IMMT2 and PRM.
After that, IMMT1 will be used for matching PRC and Xarm optical axes.

Today, after usual initial alignment, I engaged ADSs for PRX and found that my guardian code had a bug and ADS are not  engaged properly.
That seems the reason of bad alignment of PRMI yesterday. 
After fixing the bug, PRC alignment becomes good and PRMI can be locked with 3F signals (I used PRM and IMMT2 for the alignment as I mentioned above).

Figure 1 and 2 show the OLTFs of MICH and PRCL loop, respectively.
MICH UGF is about 4 Hz and phase margin is 25 degrees, which is close to the design (UGF is 4.5Hz and phase margin is 29.4 degrees).
PRCL UGF is about 20 Hz and phase margin is 47 degrees, which is also close to the design (UGF is 21Hz and phase margin is 45.6 degrees)
Since I don't have enoigh time, I cannot check how long PRMI can be locked but at least it was stably locked for 30 minutes.

Figure 3 and 4 show the beam image on POP and AS gigE camera when PRMI was locked, respectively.
Beam was not separated: PRMI seems to be locked with TEM00.
According to the AS image, the alignment seems slightly bad in pitch, which is probably due to the current initial alignment procedure (PRM are misaligned in pitch, currently).
So, I will try to engage IMMT1 ADS for more fine alignment.

To keep the good oplev values of IMMT2 for FPMI lock, I didn't record the good values by using the script.
Good alignment of IMMT2 for PRMI can be seen in fig5.

Images attached to this report
Search Help